OpenStudy (anonymous):

a grocer marks up cereal box by 20%, if a box of raisins bran costs the grocer $1.80 what is the grocers selling price

OpenStudy (anonymous):

what u send me errased

hero (hero):

Just ignore that as me playing with numbers but going wrong

hero (hero):

1.80 + .20(1.80) = 1 + 8/10 + 2/10(1 + 8/10) = 1 + 8/10 + 2/10 + 16/100 = 10/10 + 8/10 + 2/10 + 16/100 = 100/100 + 80/100 + 20/100 + 16/100 = (100 + 80 + 20 + 16)/100 = (100 + 100 + 16)/100 = (200 + 16)/100 = 216/100 = 2.16

hero (hero):

If you can decipher all of that, you're good

OpenStudy (anonymous):

i dont understand it because of the ///////

hero (hero):

It means fraction

OpenStudy (anonymous):

so how can i ge my answer

hero (hero):

\[1.80 + .20(1.80) \\= 1 + \frac{8}{10} + \frac{2}{10}(1 + \frac{8}{10}) \\= 1 + \frac{8}{10} + \frac{2}{10} + \frac{16}{100} \\= \frac{10}{10} + \frac{8}{10} + \frac{2}{10} + \frac{16}{10}\\ = \frac{100}{100} + \frac{80}{100} + \frac{20}{100} + \frac{16}{100} \\= \frac{100 + 80 + 20 + 16}{100} \\= \frac{100 + 100 + 16}{100} \\= \frac{200 + 16}{100} \\= \frac{216}{100} \\ \space \\= 2.16\]

hero (hero):

That's a long-winded solution